Illustrate with examples why “classification” and “seriation” are pre number concepts. Develop a series of three distinct activities at different levels of difficulty, to assess how for children have ability to perform any one of these two processes. Explain how these activities are at different levels of ability.



1
Expert's answer
2021-05-13T08:32:25-0400

Classification refers to sorting and grouping items by their similarities, for example, sorting toys by their colors. In contrast, seriation refers to grouping objects and items in size, position, or even location, such as sorting of pens by their length.


Moderate level:

1) 5, 7, 11, 37, 143, ?

Ans) 5 × 1 + 2 = 7

7 × 2 – 3 = 11

11 × 3 + 4 = 37

37 × 4 – 5 = 143

143 × 5 + 6 = 721


Hard level:

2)17.5, 31, 58, 98.5, 152.5, ?

Ans) 17.5 

17.5 + 1 × 13.5 = 31 

31 + 2 × 13.5 = 58 

58 + 3 × 13.5 = 98.5 

98.5 + 4 × 13.5 = 152.5 

152.5 + 5 × 13.5 = 220 


Extreme Hard level:

3) 4, 11, 25, 53,109, ?

Ans) 4 +1 × 7 = 11

11 + 2 × 7 = 25

25 + 4 × 7 = 53

53 + 8 × 7 = 109

109 + 16 × 7 = 109 + 112 = 221